OpenStudy (anonymous):

Find the product of (4x + 3y)(4x − 3y). 16x2 − 24xy + 9y2 16x2 − 9y2 16x2 + 24xy + 9y2 16x2 + 9y2

OpenStudy (anonymous):

(4x + 3y)(4x − 3y)=4x*4x-4x*3y+3y*4x-3y*3y=16x^2-12xy+12xy-9y^2=16x^2-9y^2

OpenStudy (anonymous):

b is the answer

OpenStudy (anonymous):

thx.